A PARTICLE OF MASS M=3KG MOVES ALONG A ST.LINE 4y-3x=2 with constant velocity of 5m/s.The magnitude of angular momentum about origin is...........

Grade:12

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

You have to find perpendicular distance from orgin.

It will be 2/5.

Angular mom=mvr=3*5*2/5=6kgm2/s

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

vikas askiitian expert
13 years ago

particle is moving along line 4y-3x = 2

perpendicular distance of line from origin = 4(0)-3(0)-2/(32+42)=2/5 units          (using property of line)

velocity of particle is 5m/s

angular momentam = mvrp                   (rp = perpendicular distance from line of motion to the point)

                             =3*5*2/5 = 6 units